OpenStudy (anonymous):

My directions say to solve the exponential equation. 81^7x = 9

OpenStudy (anonymous):

write 81 as 3^4 and 9 as 3^2 then solve it

OpenStudy (phi):

When you see these types of problems, look to see if you can make both sides have the same "base" 9 in this case. notice that 9^2 = 81 so you can rewrite this as \[ (9^2)^{7x} = 9^1 \]

OpenStudy (phi):

use the rule \[ (x^y)^z = x^{yz} \] so you get \[ 9^{14x} = 9^1 \]

OpenStudy (phi):

this means 14x = 1 solve for x

OpenStudy (anonymous):

Thank you.